OpenStudy (anonymous):

Select the correct graph of -x - 2y = 6.

OpenStudy (anonymous):

\[-2y=6-x\] so \[y=1/2x-3\]

OpenStudy (anonymous):

make sure of just tow points : (0,-3) and (2,-2)

OpenStudy (anonymous):

-x - 2y = 6 -2y = x + 6 y = -1/2x - 3 slope is -1/2 y intercept is (0,-3) x intercept is (-6,0) plot your intercepts (0,-3) and (-6,0) and connect them with a straight line. And you can start at your x axis (-6,0) and since slope is -1/2, come down 1 (because it is negative) and to the right 2 (because it is positive), then down 1, and to the right 2, and you should cross the y axis at (0,-3)

OpenStudy (anonymous):

sorry \[y=-1/2x-3\] (0,-3) and (-2,-2)